OpenStudy (anonymous):

I need to setup the integral in three different ways.. solid bounded between the planes x=1, y=0, and 2x+y=6, bounded below by z=0 and bounded above by z=sqrt(x). Its just setting up the different orders of integrals, no solving.

OpenStudy (watchmath):



OpenStudy (anonymous):

\[\int\limits_{1}^{3}\int\limits_{0}^{6-2x}\int\limits_{0}^{\sqrt{x} } f(x, y, z) dz dx dy\] Is this at all correct?

OpenStudy (watchmath):

you want triple or double integral?

OpenStudy (anonymous):

triple

OpenStudy (anonymous):

Sorry for the confusion.

OpenStudy (watchmath):

sorry I am in the middle of something. I can't help you right now. I'll be back if you patient :).

OpenStudy (anonymous):

Thanks!

OpenStudy (watchmath):

it is almost right but the order should be dz dy dx and also f(x,y,z)=1

OpenStudy (anonymous):

so would it be correct if it was dz dy dx? I need to find 2 more orders that's not the same, but will ultimately bring me to the same answer. I don't need to solve for f(x,y,z), just change the orders around..

OpenStudy (watchmath):

the next easiest is in the order of dz dx dy

OpenStudy (anonymous):

\[\int\limits_{0}^{6}\int\limits_{1}^{(6-y)/2}\int\limits_{0}^{\sqrt{x}} f(x,y,z) dz dx dy\] Would that be correct?

OpenStudy (watchmath):

what is the intersection between x=1 and 2x+y=6?